OpenStudy (anonymous):

-2x-y=-14 and 3x-y=11 by substitution

OpenStudy (anonymous):

Solve one equation and subsitute it into the other one.

OpenStudy (anonymous):

Which one do you want to solve?

OpenStudy (anonymous):

- 2x - y = - 14 3x - y = 11 ---> - y = - 3x + 11 ---> (-1)-y = - 3x + 11 ---> y = 3x - 11 (I multiplied this by a (-1) to make the y positive) Now we know that y = 3x - 11 and we can substitute it into the 1st equation.. -2x - y = - 14 -2x - (3x - 11) = - 14 - 2x - 3x + 11 = - 14 - 5x + 11 = - 14 - 5x = - 14 - 11 - 5x = - 25 x = 5 now substitute 5 for x into 2nd equation... y = 3x - 11 y = 3(5) - 11 y = 15 - 11 y = 4 now that we know x = 5 and y = 4 we can check by using any one of the equations.... -2x - y = - 14 -2(5) - 4 = - 14 - 10 - 4 = - 14 - 14 = - 14 (correct) 3x - y = 11 3(5) - 4 = 11 15 - 4 = 11 11 = 11 (correct) ANSWER : x = 5 and y = 4